OpenStudy (anonymous):

Help me find Q please!! For mutually exclusive events R1, R2, and R3, we have P(R1)=0.05, P(R2)=0.4, P(R3)=0.55. Also P(Q|R1)=0.5, P(Q|R2)=0.7, and P(Q|R3)=0.4. Find P(R2|Q).

OpenStudy (xapproachesinfinity):

what did you do? where is the other post hehe

OpenStudy (anonymous):

I reposted, i really need to find it haha but i know what to do once i get Q

OpenStudy (xapproachesinfinity):

So you didn't find what i told you hehe

OpenStudy (anonymous):

Im trying to! haha I've been searching everywhere

OpenStudy (xapproachesinfinity):

Okay! well unfortunately i forgot lot of stuff hehe

OpenStudy (anonymous):

ahh found it i think @xapproachesinfinity

OpenStudy (xapproachesinfinity):

Okay write down here and see how can we do it

OpenStudy (anonymous):

Its P(R2|Q)= P(Q|R2)/(P(Q|R1)*R1+P(Q|R2)*R2+P(Q|R3)*R3)

OpenStudy (xapproachesinfinity):

Are you sure lol that doesn't look good to me

OpenStudy (xapproachesinfinity):

what's the original thing you found

OpenStudy (xapproachesinfinity):

let the question open hehe

OpenStudy (anonymous):

haha whoops i closed it but i found it on my notes for my online class! i was hoping it's right

OpenStudy (xapproachesinfinity):

you are taking this course online? not quite good

OpenStudy (xapproachesinfinity):

P(R2/Q)=P(Q/R2)P(R2)/ (P(Q/R2)P(R2)+P(Q/ not R2)P( not R2) i guess this the formula

OpenStudy (xapproachesinfinity):

1-P(R2)=P( not R2)=1-0.4=0.6

OpenStudy (anonymous):

what do you mean not R2?

OpenStudy (xapproachesinfinity):

Complement of R2 is not R2 it mean R2 not accuring

OpenStudy (xapproachesinfinity):

they usually denote with a c on top if you check your text book you will find it that way

OpenStudy (xapproachesinfinity):

let's event A is it is raining today not A will be it is not raining do you get it?

OpenStudy (xapproachesinfinity):

i think this is now simple!

OpenStudy (xapproachesinfinity):

we already got P(Q/R2)=0.7

OpenStudy (xapproachesinfinity):

we know P(R2)=0.4

OpenStudy (xapproachesinfinity):

we are missing one thing P(Q/not R2)

OpenStudy (xapproachesinfinity):

P(Q/not R2)=P(Q and not R2)/P(R2) P(Q and not R2)=P(Q)P(not R2)

OpenStudy (anonymous):

but we still dont know P(Q)?

OpenStudy (xapproachesinfinity):

Darn it! again P(Q) is missing now that i think about it there is a reason why they give you all those probab so you can use it to get that P(Q)

OpenStudy (xapproachesinfinity):

I would like to play with this little more but im running out of time here

OpenStudy (xapproachesinfinity):

@jdoe0001

OpenStudy (xapproachesinfinity):

Note that P(Q/R3)=P(R2)

OpenStudy (xapproachesinfinity):

So P(Q/R3)=P(Q)P(R3)/P(R3)=P(R2) hence P(Q)=P(R2)=0.4

OpenStudy (xapproachesinfinity):

is what you wrote correct! make sure

OpenStudy (anonymous):

Ill double check I have to go run an errand but thank you so much!

OpenStudy (xapproachesinfinity):

we almost done! just plug the numbers! and you are welcome